# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 5.4. Products of $\mathrm{GL}_{n}$’s
The requirements that $\chi$ is supersingular and $S=S_{\xi}$ ensure that there exist $s_{i},s_{i}^{\prime}\in S_{i}$, adjacent in the affine Dynkin diagram, such that $\chi(T_{\hat{s_{i}}})\neq\chi(T_{\hat{s_{i}^{\prime}}})$. If there is $d>2$ dividing $n_{i}$ then we automatically have $n_{i}>2$ and we may define $F\subseteq\overline{C}$ such that $S_{F}$ is equal to the orbit of $\{s_{i},s_{i}^{\prime}\}$ under the subgroup $\langle\omega_{i}^{d}\rangle$. By construction we have $\Omega_{F}=\langle\omega_{i}^{d}\rangle\times\Omega_{T^{\prime}}\times\prod_{j\neq i}\Omega_{j}$, and 5.10(ii) gives that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective as required. ∎
The above will not help in the event that $d_{i}=2$ for all $i$. We deal with that case directly now, first introducing some notation. For each $1\leq i\leq r$ we let $[r]_{i}:=\{1\leq j\leq r\mid j\neq i\}$, and for any $\mathbf{j}\subseteq[r]_{i}$ we let $\hat{\omega}_{\mathbf{j}}:=\prod_{j\in\mathbf{j}}\hat{\omega}_{j}\in\widetilde{\Omega}$. We also fix a generating set $\omega_{T^{\prime},1},\ldots,\omega_{T^{\prime},l}$ for $\Omega_{T^{\prime}}$ and write $\hat{\omega}_{T^{\prime},1},\ldots,\hat{\omega}_{T^{\prime},l}$ for the corresponding lifts to $\widetilde{\Omega}_{T^{\prime}}$.
@@
For each $1\leq i\leq r$, we write $\lambda_{i}$ and $\lambda^{\prime}_{i}$ for the scalar by which $\hat{\omega}_{i}^{2}$ acts on $V$ and $V^{\prime}$, respectively. Similarly, for each $1\leq j\leq l$, we write $\nu_{j}$ and $\nu_{j}^{\prime}$ for the scalar by which $\hat{\omega}_{T^{\prime},j}$ acts on $V$ and $V^{\prime}$ respectively. From our assumption on $\chi$ and by 5.5(iii), there is a face $F\subseteq\overline{C}$ for which $\chi|_{{\mathcal{H}}_{F}}$ is not projective. Note that we have $\widetilde{\Omega}_{T^{\prime}}\subseteq\widetilde{\Omega}_{F,\chi}$ and $\hat{\omega}_{i}^{2}\in\widetilde{\Omega}_{F,\chi}$ for each $i$ such that $n_{i}=2$. Applying 5.23 we deduce that $\nu_{j}=\nu_{j}^{\prime}$ for all $j$ and that $\lambda_{i}=\lambda^{\prime}_{i}$ whenever $n_{i}=2$.
@@
We are thus left to show that $\lambda_{i}=\lambda^{\prime}_{i}$ for all $i$ for which $n_{i}>2$, i.e. $m_{i}>1$. Pick $s_{0}\in S_{i}$ and write $s_{j}=\omega_{i}^{-j}s_{0}\omega_{i}^{j}$ for $11\leq j\leq n_{i}-$ 1. By our assumptions on $\chi$, without loss of generality we may assume that $J\cap S_{i}=\{s_{2j}\mid 0\leq j\leq m_{i}-1\}$. Similarly $\chi^{\hat{\omega}_{i}}=(J^{\omega_{i}},\xi)$ then satisfies $J^{\omega_{i}}\cap S_{i}=\{s_{2j+1}\mid 0\leq j\leq m_{i}-1\}$. Note that any $\widetilde{\Omega}$-conjugate of $\chi$ is non-projective when restricted to any ${\mathcal{H}}_{F}$ such that $S_{F}\cap S_{i}$ contains two adjacent simple affine reflections in the affine Dynkin diagram, c.f. 5.10(ii). Fix an isomorphism $f:{\mathfrak{m}}\to{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$, write $V=kv$ and $V^{\prime}=kv^{\prime}$, and for each $F\subseteq\overline{C}$ let $f_{F}:{\mathfrak{m}}\to{\mathfrak{m}}^{\prime}$ be a lift in ${\operatorname{Mod}}({\mathcal{H}}_{F})$ of ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{F}})(f)$.
@@
Pick a prime factor $\ell$ of $m_{i}$. Similarly to the proof of 5.29, we let $F$ be the face given by taking $S_{F}$ to be the $\langle\omega_{i}^{2\ell}\rangle$-orbit of $\{s_{0},s_{1}\}$ so that $\chi|_{{\mathcal{H}}_{F}}$ and $(\chi^{\hat{\omega}_{i}})|_{{\mathcal{H}}_{F}}$ are not projective. We also have that $\chi|_{{\mathcal{H}}_{F\omega_{i}^{-1}}}$ and $(\chi^{\hat{\omega}})|_{{\mathcal{H}}_{F\omega_{i}^{-1}}}$ are not projective.
@@
Next, since $s_{0}\in S_{F}\cap S_{F\omega_{i}^{-1}}$ it follows that $\left|S_{F}\cup S_{F\omega_{i}^{-1}}\right|<2\cdot\left|S_{F}\right|=4m_{i}/\ell\leq n_{i}$ and hence there is a face $F^{\prime}\subseteq\overline{F}$ with $S_{F^{\prime}}=S_{F}\cup S_{F\omega_{i}^{-1}}$. Again both $\chi$ and $\chi^{\hat{\omega}_{i}}$ have non-projective restrictions to ${\mathcal{H}}_{F^{\prime}}$. Also note that for all $\tilde{\omega}\in\widetilde{\Omega}$ we have that $\chi^{\tilde{\omega}}|_{{\mathcal{H}}_{F^{\prime}}}$ is equal to either $\chi|_{{\mathcal{H}}_{F^{\prime}}}$ or $(\chi^{\hat{\omega}_{i}})|_{{\mathcal{H}}_{F^{\prime}}}$. By the proof of 5.23 and 5.6 we deduce that, with respect to the bases $\bigcup_{\mathbf{j}\subseteq[r]_{i}}\{v\otimes T_{\hat{\omega}_{\mathbf{j}}},v\otimes T_{\hat{\omega}_{\mathbf{j}}}T_{\hat{\omega}_{i}}\}$ and $\bigcup_{\mathbf{j}\subseteq[r]_{i}}\{v^{\prime}\otimes T_{\hat{\omega}_{\mathbf{j}}},v^{\prime}\otimes T_{\hat{\omega}_{\mathbf{j}}}T_{\hat{\omega}_{i}}\}$ of ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$ respectively, $f_{F^{\prime}}$ is a block diagonal matrix $A$ whose blocks are $2\times 2$-matrices, indexed by the subsets $\mathbf{j}\subseteq[r]_{i}$, equal to $\begin{pmatrix}\mu_{\mathbf{j}}&0\\ 0&\mu^{\prime}_{\mathbf{j}}\end{pmatrix}$ for some scalars $\mu_{\mathbf{j}},\mu^{\prime}_{\mathbf{j}}\in k^{\times}$. Similarly, $f_{F}$ and $f_{F\omega_{i}^{-1}}$ are also given by block diagonal matrices of the same form as above, and 5.6 ensures that both of these are actually equal to $A$ since $f_{F}\sim(f_{F^{\prime}})|_{{\mathcal{H}}_{F}}$ and $f_{F\omega_{i}^{-1}}\sim(f_{F^{\prime}})|_{{\mathcal{H}}_{F\omega_{i}^{-1}}}$.
@@
Now, with respect to the above bases, the action of $T_{\hat{\omega}_{i}}$ on ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$ is given by two block diagonal matrices $B$ and $B^{\prime}$ respectively, whose $2\times 2$-blocks are indexed by the subsets $\mathbf{j}\subseteq[r]_{i}$ as above but with each block equal to $\begin{pmatrix}0&1\\ \lambda_{i}&0\end{pmatrix}$ and $\begin{pmatrix}0&1\\ \lambda_{i}^{\prime}&0\end{pmatrix}$ respectively (where we view these matrices as acting on the right). Using that $f_{F}=f_{F\omega_{i}^{-1}}\cdot\omega$, c.f. 5.21 and 5.6, we deduce that $A=B^{-1}AB^{\prime}$. Looking at each $2\times 2$-block, this tells us that
@@
Assuming that $(n_{1},\ldots,n_{r})\neq(2,\ldots,2)$, note that 5.30 is the only situation where we could not appeal to 5.24. Indeed, by the description of $\chi$ as given in the proof we obtain from 5.10(ii) that a face $F\subseteq\overline{C}$ satisfies that $\chi|_{{\mathcal{H}}_{F}}$ is not projective if and only if $S_{F}$ contains two adjacent nodes in the affine Dynkin diagram. If those belong to $S_{i}$, say, then $F\omega_{i}^{2}\neq F$ and we deduce that $\widetilde{\Omega}_{\chi}\nsubseteq\widetilde{\Omega}_{F}$.
@@
Suppose that $G=\mathrm{GL}_{n_{1}}\times\cdots\times\mathrm{GL}_{n_{r}}\times\mathbb{T}^{\prime}$, where $n_{1}\geq\ldots\geq n_{r}\geq 2$ and $\mathbb{T}^{\prime}\cong\mathbb{G}_{m}^{l}$ is a split torus. Let ${\mathfrak{m}}=(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ and ${\mathfrak{m}}^{\prime}={\mathcal{H}}\otimes_{{\mathcal{H}}_{\chi^{\prime}}}(\chi^{\prime}\otimes V^{\prime})$ be two simple supersingular ${\mathcal{H}}$-modules of infinite projective dimension over ${\mathcal{H}}$, where $\chi=(J,\xi)$ and $\chi^{\prime}=(J^{\prime},\xi^{\prime})$.
@@
If $S\neq S_{\xi}$ then 5.27 implies our claim, so that we only need to consider the case $S=S_{\xi}$. If $d_{i}>2$ for some $1\leq i\leq r$ then 5.29 and 5.24 together imply that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}({\mathcal{H}})$. If $d_{i}=2$ for all $1\leq i\leq r$, we get ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}({\mathcal{H}})$ from 5.30. This concludes the proof of (5.12).
@@
When $(n_{1},n_{2},\ldots,n_{r})\neq(3,2,\ldots,2)$, 5.17 implies that $\chi$ and $\chi^{\prime}$ must be $\widetilde{\Omega}$-conjugate for ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$ to be isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$. Thus (i) follows immediately from (5.12). We are therefore left to consider the case where $(n_{1},n_{2},\ldots,n_{r})=(3,2,\ldots,2)$ and $\chi,\chi^{\prime}$ are not $\widetilde{\Omega}$-conjugate. Then 5.17 again implies that, up to permuting ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$, we must have $\chi$ and $\chi^{\prime}$ as in the statement in order for ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$ to be isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$. Note that we then have $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\prime}}=\widetilde{\Omega}_{T^{\prime}}\rtimes(\langle\hat{\omega}_{1}^{3}\rangle\times\langle\hat{\omega}_{2}^{2}\rangle\times\cdots\times\langle\hat{\omega}_{r}^{2}\rangle)$.
@@
We must finally show that under these assumptions on $\chi$ and $\chi^{\prime}$, ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ if and only if $\lambda_{i}=\lambda_{i}^{\prime}$ and $\nu_{j}=\nu_{j}^{\prime}$ for all $1\leq i\leq r$ and all $1\leq j\leq l$. Assume first that these equalities all hold. We argue similarly to the proof of 5.7. By the product decomposition of $\mathbb{G}$, we have an algebra isomorphism ${\mathcal{H}}\cong{\mathcal{H}}_{1}\otimes_{k}{\mathcal{H}}_{\geq 2}$ where ${\mathcal{H}}_{1}$ is the pro-$p$ Iwahori-Hecke algebra of $\mathrm{GL}_{3}(\mathfrak{F})$ and ${\mathcal{H}}_{\geq 2}$ is the pro-$p$ Iwahori-Hecke algebra of $\mathrm{GL}_{2}(\mathfrak{F})\times\cdots\times\mathrm{GL}_{2}(\mathfrak{F})\times T^{\prime}$ (with $r-1$ $\mathrm{GL}_{2}$-factors). There is a similar decomposition ${\mathcal{H}_{\operatorname{aff}}}={\mathcal{H}}_{{\operatorname{aff}},1}\otimes_{k}{\mathcal{H}}_{{\operatorname{aff}},\geq 2}$ and correspondingly we have $\chi=\chi_{1}\otimes_{k}\chi_{\geq 2}$. Also, we have a decomposition $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi,1}\times\widetilde{\Omega}_{\chi,\geq 2}$, and we may thus write $V=V_{1}\otimes_{k}V_{\geq 2}$ with $V_{1}=kv_{1}$ and $V_{\geq 2}=kv_{\geq 2}$ one-dimensional representations of $\widetilde{\Omega}_{\chi,1}$ and $\widetilde{\Omega}_{\chi,\geq 2}$ respectively.
@@
Write $v=v_{1}\otimes v_{\geq 2}\in V$. By considering the basis $\bigcup_{\mathbf{j}\subseteq[r]_{1}}\{v\otimes T_{\hat{\omega}_{\mathbf{j}}},v\otimes T_{\hat{\omega}_{\mathbf{j}}}T_{\hat{\omega}_{1}},v\otimes T_{\hat{\omega}_{\mathbf{j}}}T_{\hat{\omega}^{2}_{1}}\}$ of ${\mathfrak{m}}$, we see by inspection that ${\mathfrak{m}}\cong{\mathfrak{m}}_{1}\otimes_{k}{\mathfrak{m}}_{\geq 2}$ where ${\mathfrak{m}}_{1}\cong(\chi_{1}\otimes V_{1})\otimes_{{\mathcal{H}}_{\chi,1}}{\mathcal{H}}_{1}$ and ${\mathfrak{m}}_{\geq 2}\cong(\chi_{\geq 2}\otimes V_{\geq 2})\otimes_{{\mathcal{H}}_{\chi,\geq 2}}{\mathcal{H}}_{\geq 2}$. We note that these are simple supersingular modules of ${\mathcal{H}}_{1}$ and ${\mathcal{H}}_{\geq 2}$ respectively, and that ${\operatorname{pd}}_{{\mathcal{H}}_{\geq 2}}({\mathfrak{m}}_{\geq 2})<\infty$ by 5.4. Completely analogously, we have a decomposition ${\mathfrak{m}}^{\prime}\cong{\mathfrak{m}}^{\prime}_{1}\otimes_{k}{\mathfrak{m}}^{\prime}_{\geq 2}$ where the factors are simple supersingular modules over ${\mathcal{H}}_{1}$ and ${\mathcal{H}}_{\geq 2}$ respectively. Furthermore, by our assumptions on $\chi$ and $\chi^{\prime}$ we have ${\mathfrak{m}}_{\geq 2}\cong{\mathfrak{m}}^{\prime}_{\geq 2}$. We then deduce from 5.19 and 5.14 that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ as required.

### v4

@@ 5.4. Products of $\mathrm{GL}_{n}$’s
The requirements that $\chi$ is supersingular and $S=S_{\xi}$ ensure that there exist $s_{i},s_{i}^{\prime}\in S_{i}$, adjacent in the affine Dynkin diagram, such that $\chi(T_{\hat{s_{i}}})\neq\chi(T_{\hat{s_{i}^{\prime}}})$. If there is $d>2$ dividing $n_{i}$ then we automatically have $n_{i}>2$ and we may define $F\subseteq\overline{C}$ such that $S_{F}$ is equal to the orbit of $\{s_{i},s_{i}^{\prime}\}$ under the subgroup $\langle\omega_{i}^{d}\rangle$. By construction we have $\Omega_{F}=\langle\omega_{i}^{d}\rangle\times\Omega_{T^{\prime}}\times\prod_{j\neq i}\Omega_{j}$, and 5.12(ii) gives that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective as required. ∎
We will be unable to use the above with $d=d_{j}$ (some $j$) in the case that $d_{i}=2$ for all $1\leq i\leq r$. We deal with that case next, first introducing some notation. For each $1\leq i\leq r$ we let $[r]_{i}:=\{1\leq j\leq r\mid j\neq i\}$, and for any $\mathbf{j}\subseteq[r]_{i}$ we let $\hat{\omega}_{\mathbf{j}}:=\prod_{j\in\mathbf{j}}\hat{\omega}_{j}\in\widetilde{\Omega}$. We fix a generating set $\omega_{T^{\prime},1},\ldots,\omega_{T^{\prime},l}$ for $\Omega_{T^{\prime}}$ and write $\hat{\omega}_{T^{\prime},1},\ldots,\hat{\omega}_{T^{\prime},l}$ for the corresponding lifts to $\widetilde{\Omega}_{T^{\prime}}$.
@@
For each $1\leq i\leq r$, we write $\lambda_{i}$ and $\lambda^{\prime}_{i}$ for the scalar by which $\hat{\omega}_{i}^{2}$ acts on $V$ and $V^{\prime}$, respectively. Similarly, for each $1\leq j\leq l$, we write $\nu_{j}$ and $\nu_{j}^{\prime}$ for the scalar by which $\hat{\omega}_{T^{\prime},j}$ acts on $V$ and $V^{\prime}$ respectively. From our assumption on $\chi$ and by 5.6(iii), there is a face $F\subseteq\overline{C}$ for which $\chi|_{{\mathcal{H}}_{F}}$ is not projective. Note that we have $\widetilde{\Omega}_{T^{\prime}}\subseteq\widetilde{\Omega}_{F,\chi}$ and $\hat{\omega}_{i}^{2}\in\widetilde{\Omega}_{F,\chi}$ for each $i$ such that $n_{i}=2$. Applying 5.25 we deduce that $\nu_{j}=\nu_{j}^{\prime}$ for all $j$ and that $\lambda_{i}=\lambda^{\prime}_{i}$ whenever $n_{i}=2$.
@@
We are thus left to show that $\lambda_{i}=\lambda^{\prime}_{i}$ for all $i$ for which $n_{i}>2$, i.e. $m_{i}>1$. Pick $s_{0}\in S_{i}$ and write $s_{j}=\omega_{i}^{-j}s_{0}\omega_{i}^{j}$ for $11\leq j\leq n_{i}-$ 1. By our assumptions on $\chi$, without loss of generality we may assume that $J\cap S_{i}=\{s_{2j}\mid 0\leq j\leq m_{i}-1\}$. Similarly $\chi^{\hat{\omega}_{i}}=(J^{\omega_{i}},\xi)$ then satisfies $J^{\omega_{i}}\cap S_{i}=\{s_{2j+1}\mid 0\leq j\leq m_{i}-1\}$. Note that any $\widetilde{\Omega}$-conjugate of $\chi$ is non-projective when restricted to any ${\mathcal{H}}_{F}$ such that $S_{F}\cap S_{i}$ contains two adjacent simple affine reflections in the affine Dynkin diagram, cf. 5.12(ii). Fix an isomorphism $f:{\mathfrak{m}}\to{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$, write $V=kv$ and $V^{\prime}=kv^{\prime}$, and for each $F\subseteq\overline{C}$ let $f_{F}:{\mathfrak{m}}\to{\mathfrak{m}}^{\prime}$ be a lift in ${\operatorname{Mod}}({\mathcal{H}}_{F})$ of ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{F}})(f)$.
@@
Pick a prime factor $\ell$ of $m_{i}$. Similarly to the proof of 5.32, we let $F$ be the face given by taking $S_{F}$ to be the $\langle\omega_{i}^{2\ell}\rangle$-orbit of $\{s_{0},s_{1}\}$ so that $\chi|_{{\mathcal{H}}_{F}}$ and $(\chi^{\hat{\omega}_{i}})|_{{\mathcal{H}}_{F}}$ are not projective. We also have that $\chi|_{{\mathcal{H}}_{F\omega_{i}^{-1}}}$ and $(\chi^{\hat{\omega}})|_{{\mathcal{H}}_{F\omega_{i}^{-1}}}$ are not projective.
@@
Next, since $s_{0}\in S_{F}\cap S_{F\omega_{i}^{-1}}$ it follows that $\left|S_{F}\cup S_{F\omega_{i}^{-1}}\right|<2\cdot\left|S_{F}\right|=4m_{i}/\ell\leq n_{i}$ and hence there is a face $F^{\prime}\subseteq\overline{F}$ with $S_{F^{\prime}}=S_{F}\cup S_{F\omega_{i}^{-1}}$. Again both $\chi$ and $\chi^{\hat{\omega}_{i}}$ have non-projective restrictions to ${\mathcal{H}}_{F^{\prime}}$. Also note that for all $\tilde{\omega}\in\widetilde{\Omega}$ we have that $\chi^{\tilde{\omega}}|_{{\mathcal{H}}_{F^{\prime}}}$ is equal to either $\chi|_{{\mathcal{H}}_{F^{\prime}}}$ or $(\chi^{\hat{\omega}_{i}})|_{{\mathcal{H}}_{F^{\prime}}}$. By the proof of 5.25 and 5.7 we deduce that, with respect to the bases $\bigcup_{\mathbf{j}\subseteq[r]_{i}}\{v\otimes T_{\hat{\omega}_{\mathbf{j}}},v\otimes T_{\hat{\omega}_{\mathbf{j}}}T_{\hat{\omega}_{i}}\}$ and $\bigcup_{\mathbf{j}\subseteq[r]_{i}}\{v^{\prime}\otimes T_{\hat{\omega}_{\mathbf{j}}},v^{\prime}\otimes T_{\hat{\omega}_{\mathbf{j}}}T_{\hat{\omega}_{i}}\}$ of ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$ respectively, $f_{F^{\prime}}$ is a block diagonal matrix $A$ whose blocks are $2\times 2$-matrices, indexed by the subsets $\mathbf{j}\subseteq[r]_{i}$, equal to $\begin{pmatrix}\mu_{\mathbf{j}}&0\\ 0&\mu^{\prime}_{\mathbf{j}}\end{pmatrix}$ for some scalars $\mu_{\mathbf{j}},\mu^{\prime}_{\mathbf{j}}\in k^{\times}$. Similarly, $f_{F}$ and $f_{F\omega_{i}^{-1}}$ are also given by block diagonal matrices of the same form as above, and 5.7 ensures that both of these are actually equal to $A$ since $f_{F}\sim(f_{F^{\prime}})|_{{\mathcal{H}}_{F}}$ and $f_{F\omega_{i}^{-1}}\sim(f_{F^{\prime}})|_{{\mathcal{H}}_{F\omega_{i}^{-1}}}$.
@@
Now, with respect to the above bases, the action of $T_{\hat{\omega}_{i}}$ on ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$ is given by two block diagonal matrices $B$ and $B^{\prime}$ respectively, whose $2\times 2$-blocks are indexed by the subsets $\mathbf{j}\subseteq[r]_{i}$ as above but with each block equal to $\begin{pmatrix}0&1\\ \lambda_{i}&0\end{pmatrix}$ and $\begin{pmatrix}0&1\\ \lambda_{i}^{\prime}&0\end{pmatrix}$ respectively (where we view these matrices as acting on the right). Using that $f_{F}=f_{F\omega_{i}^{-1}}\cdot\omega$, cf. 5.24 and 5.7, we deduce that $A=B^{-1}AB^{\prime}$. Looking at each $2\times 2$-block, this tells us that
@@
Assuming that $(n_{1},\ldots,n_{r})\neq(2,\ldots,2)$, note that 5.33 is the only situation where we could not appeal to 5.26. Indeed, by the description of $\chi$ as given in the proof we obtain from 5.12(ii) that a face $F\subseteq\overline{C}$ satisfies that $\chi|_{{\mathcal{H}}_{F}}$ is not projective if and only if $S_{F}$ contains two adjacent nodes in the affine Dynkin diagram. If those belong to $S_{i}$, say, then $F\omega_{i}^{2}\neq F$ and we deduce that $\widetilde{\Omega}_{\chi}\nsubseteq\widetilde{\Omega}_{F}$.
@@
Suppose that $\mathbb{G}=\mathrm{GL}_{n_{1}}\times\cdots\times\mathrm{GL}_{n_{r}}\times\mathbb{T}^{\prime}$, where $n_{1}\geq\ldots\geq n_{r}\geq 2$ and $\mathbb{T}^{\prime}\cong\mathbb{G}_{m}^{l}$ is a split torus. Let ${\mathfrak{m}}=(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ and ${\mathfrak{m}}^{\prime}=(\chi^{\prime}\otimes V^{\prime})\otimes_{{\mathcal{H}}_{\chi^{\prime}}}{\mathcal{H}}$ be two simple supersingular ${\mathcal{H}}$-modules of infinite projective dimension over ${\mathcal{H}}$, where $\chi=(J,\xi)$ and $\chi^{\prime}=(J^{\prime},\xi^{\prime})$.
@@
If $S\neq S_{\xi}$ then 5.30 implies our claim, so that we only need to consider the case $S=S_{\xi}$. If $d_{i}>2$ for some $1\leq i\leq r$ then 5.32 and 5.26 together imply that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}({\mathcal{H}})$. If $d_{i}=2$ for all $1\leq i\leq r$, we get ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}({\mathcal{H}})$ from 5.33. This concludes the proof of (5.14).
@@
When $(n_{1},n_{2},\ldots,n_{r})\neq(3,2,\ldots,2)$, 5.19 implies that $\chi$ and $\chi^{\prime}$ must be $\widetilde{\Omega}$-conjugate for ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$ to be isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$. Thus (i) follows immediately from (5.14). We are therefore left to consider the case where $(n_{1},n_{2},\ldots,n_{r})=(3,2,\ldots,2)$ and $\chi,\chi^{\prime}$ are not $\widetilde{\Omega}$-conjugate. Then 5.19 again implies that, up to permuting ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$, we must have $\chi$ and $\chi^{\prime}$ as in the statement in order for ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$ to be isomorphic in ${\operatorname{Ho}}({\mathcal{H}})$. Note that we then have $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\prime}}=\widetilde{\Omega}_{T^{\prime}}\rtimes(\langle\hat{\omega}_{1}^{3}\rangle\times\langle\hat{\omega}_{2}^{2}\rangle\times\cdots\times\langle\hat{\omega}_{r}^{2}\rangle)$.
@@
We must finally show that under these assumptions on $\chi$ and $\chi^{\prime}$, ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ if and only if $\lambda_{i}=\lambda_{i}^{\prime}$ and $\nu_{j}=\nu_{j}^{\prime}$ for all $1\leq i\leq r$ and all $1\leq j\leq l$. Assume first that these equalities all hold. We argue similarly to the proof of 5.8. By the product decomposition of $\mathbb{G}$, we have an algebra isomorphism ${\mathcal{H}}\cong{\mathcal{H}}_{1}\otimes_{k}{\mathcal{H}}_{\geq 2}$ where ${\mathcal{H}}_{1}$ is the pro-$p$ Iwahori-Hecke algebra of $\mathrm{GL}_{3}(\mathfrak{F})$ and ${\mathcal{H}}_{\geq 2}$ is the pro-$p$ Iwahori-Hecke algebra of $\mathrm{GL}_{2}(\mathfrak{F})\times\cdots\times\mathrm{GL}_{2}(\mathfrak{F})\times T^{\prime}$ (with $r-1$ $\mathrm{GL}_{2}$-factors). There is a similar decomposition ${\mathcal{H}_{\operatorname{aff}}}={\mathcal{H}}_{{\operatorname{aff}},1}\otimes_{k}{\mathcal{H}}_{{\operatorname{aff}},\geq 2}$ and correspondingly we have $\chi=\chi_{1}\otimes_{k}\chi_{\geq 2}$. Also, we have a decomposition $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi,1}\times\widetilde{\Omega}_{\chi,\geq 2}$, and we may thus write $V=V_{1}\otimes_{k}V_{\geq 2}$ with $V_{1}=kv_{1}$ and $V_{\geq 2}=kv_{\geq 2}$ one-dimensional representations of $\widetilde{\Omega}_{\chi,1}$ and $\widetilde{\Omega}_{\chi,\geq 2}$ respectively.
@@
Write $v=v_{1}\otimes v_{\geq 2}\in V$. By considering the basis $\bigcup_{\mathbf{j}\subseteq[r]_{1}}\{v\otimes T_{\hat{\omega}_{\mathbf{j}}},v\otimes T_{\hat{\omega}_{\mathbf{j}}}T_{\hat{\omega}_{1}},v\otimes T_{\hat{\omega}_{\mathbf{j}}}T_{\hat{\omega}^{2}_{1}}\}$ of ${\mathfrak{m}}$, we see by inspection that ${\mathfrak{m}}\cong{\mathfrak{m}}_{1}\otimes_{k}{\mathfrak{m}}_{\geq 2}$ where ${\mathfrak{m}}_{1}\cong(\chi_{1}\otimes V_{1})\otimes_{{\mathcal{H}}_{\chi,1}}{\mathcal{H}}_{1}$ and ${\mathfrak{m}}_{\geq 2}\cong(\chi_{\geq 2}\otimes V_{\geq 2})\otimes_{{\mathcal{H}}_{\chi,\geq 2}}{\mathcal{H}}_{\geq 2}$. We note that these are simple supersingular modules of ${\mathcal{H}}_{1}$ and ${\mathcal{H}}_{\geq 2}$ respectively, and that ${\operatorname{pd}}_{{\mathcal{H}}_{\geq 2}}({\mathfrak{m}}_{\geq 2})<\infty$ by 5.4. Completely analogously, we have a decomposition ${\mathfrak{m}}^{\prime}\cong{\mathfrak{m}}^{\prime}_{1}\otimes_{k}{\mathfrak{m}}^{\prime}_{\geq 2}$ where the factors are simple supersingular modules over ${\mathcal{H}}_{1}$ and ${\mathcal{H}}_{\geq 2}$ respectively. Furthermore, by our assumptions on $\chi$ and $\chi^{\prime}$ we have ${\mathfrak{m}}_{\geq 2}\cong{\mathfrak{m}}^{\prime}_{\geq 2}$. We then deduce from 5.21 and 5.16 that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Ho}}({\mathcal{H}})$ as required.